γ-ray photons with a wavelength of 2.41×10−5 nm .

Express the energy numerically in kilojoules per mole
Answer #1

1=2.41x 10-5 nm lam= 10-9m d=2.41x10-14m E = 6.626X10-34 x 3 x 10 2.41610-14 E = 6.626 X 10-34+8-(-14) 2.41 t= 6.626x3 x 10-3